OpenStudy (anonymous):

Which equation is equivalent to 2x + 6y = 12?

OpenStudy (anonymous):

I am assuming they want it in point slope form. 2x + 6y = 12 Isolate y. 6y = -2x + 12 y = -2/6x + 2